13) Determine the decision criterion for rejecting the null hypothesis in the given hypothesis test; i.e., describe the values of the test statistic that would result in rejection of the null hypothesis.
Suppose you wish to test the claim that μd, the mean value of the differences d for a population of paired data, is greater than 0. Given a sample of n = 15 and a significance level of α = 0.01, what criterion would be used for rejecting the null hypothesis?   
Answer #1

13) Determine the decision criterion for rejecting the null hypothesis in the given hypothesis test; i.e., describe the values of the test statistic that would result in rejection of the null hypothesis.
Suppose you wish to test the claim that μd, the mean value of the differences d for a population of paired data, is greater than 0. Given a sample of n = 15 and a significance level of α = 0.01, what criterion would be used for rejecting the null hypothesis?   

Degrees of freedom = n-1= 15-1=14

The claim is μd > 0, this is upper tail test

Table value of t at 14 = 2.624 ( upper tail value at 0.01 level)

If the calculated t value > 2.624 reject the null hypothesis.
Answer #2

The 24 consonant sounds comprise six stops (plosives): p, b, t, d, k, g; the fricatives f, v, θ (as in thin), ð [eth] (as in then), s, z, [esh] (as in ship), Ʒ (as in pleasure), and h; two affricatives: t∫ (as in church) and dƷ (as the j in jam); the nasals m, n, ŋ (the sound that occurs at the end of words such as young); the lateral l; the postalveolar or retroflex r; and the semivowels j (often spelled y) and w. These remain fairly stable, but Inland Northern American differs from RP in two respects: (1) r following vowels is preserved in words such as door, flower, and harmony, whereas it is lost in RP; (2) t between vowels is voiced, so that metal and matter sound very much like British medal and madder, although the pronunciation of this t is softer and less aspirated, or breathy, than the d of British English.

answered by: Gamesome Goshawk